OpenStudy (anonymous):

@HelpMePlease Find the first,Fourth, and Eighth terms of the sequence A(n)=-3*2 ^n-1 a. -6 -48 -768 b -3 -24 -384 c 1 -216 -276 -279,936 d -12 -96 -1,536

OpenStudy (freckles):

first term can be found by replacing n with 1 4th term can be found by replacing n with 4 8th term can be found by replacing n with 8

OpenStudy (anonymous):

yA

OpenStudy (anonymous):

I tryed that i cant find the answer

OpenStudy (freckles):

\[A(n)=-3 \cdot 2^n-1\] replace n with 1 \[A(1)=-3 \cdot 2^1-1 \\ A(1)=-3 \cdot 2-1 \\ A(1)=-6-1 \\ A(1)=-7\] I see the problem now... -7 isn't an option for the 1st term

OpenStudy (freckles):

Or you sure you entered in the correct expression for A(n)?

OpenStudy (freckles):

maybe you meant A(n)=-3*2^(n-1)?

OpenStudy (anonymous):

thats what i said bro

OpenStudy (freckles):

if so then you would have this \[A(n)=-3 \cdot 2^{n-1}\]

OpenStudy (freckles):

no it isn't you had -3*2^n-1

OpenStudy (freckles):

which means \[A(n)=-3 \cdot 2^n-1\]

OpenStudy (freckles):

anyways if you mean A(n)=-3*2^(n-1) \[A(n)=-3 \cdot 2^{n-1}\]

OpenStudy (freckles):

replace n with 1 can you tell me what you get as the output

OpenStudy (anonymous):

@xo_kansasprincess_xo

OpenStudy (freckles):

keep in mind a^0=1 when a doesn't equal 0

OpenStudy (freckles):

let me know when you have evaluated \[A(1)=-3 \cdot 2^{1-1}\] and I will check it for you

OpenStudy (anonymous):

@Poppy_leelee14

OpenStudy (anonymous):

Mathway.com

OpenStudy (freckles):

if you need help evaluating it can you please tell me where you are having trouble?

OpenStudy (freckles):

like 1-1 =?

OpenStudy (freckles):

when you have 1 apple and you decide to eat that apple you have how many apples left?

OpenStudy (anonymous):

so what would the answer be

OpenStudy (freckles):

have you evaluated \[-3 \cdot 2^{1-1} \text{ yet }\]

OpenStudy (freckles):

I was trying to help you but you haven't responded to me

OpenStudy (freckles):

first do the 1-1

OpenStudy (anonymous):

yah cause i dont get it and i dont want to look like a ideat

OpenStudy (freckles):

you can do 1-1 though

OpenStudy (freckles):

like if you have exactly 1 dollar and you give that dollar away then how many dollars do you have

OpenStudy (anonymous):

1-1=0

OpenStudy (freckles):

yes you did part of it! :) \[A(1)=-3 \cdot 2^{1-1}=-3 \cdot 2^0\] and I told you what a^0 was for any number a that isn't 0

OpenStudy (freckles):

what is 2^0 ?

OpenStudy (anonymous):

would u mind helping me on google please

OpenStudy (anonymous):

@freckles did u get my pm

OpenStudy (freckles):

I think it is easier for me to use OpenStudy

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

i just did not want to have every one see how dum i am

OpenStudy (freckles):

No one thinks you are dumb.

OpenStudy (freckles):

This is a learning site. That is all anyone expects anyone to be doing here.

OpenStudy (anonymous):

i do because people come to me for help and i am in need of help